OpenStudy (anonymous):

The vertex of y = 9 - 8x - x2 is (-4, 25) True False

OpenStudy (anonymous):

True

OpenStudy (mertsj):

\[x=\frac{8}{-2}=-4\] y=9-8(-4)-(-4)^2=9+32-16=25

OpenStudy (anonymous):

If you are not convinced by the good solution of @Mertsj , see the attached graph.